November 1996  ICTCM IX (Reno)  32

Advertisements

• Using Maple to Explore Mathematical Models in Biology Matt Miller (USC), Friday, 9am, Crystal 1

• Computer Algebra and the WWW in the Math Curriculum Murphy, Miller, Meade, Friday, 12pm, Crystal 2

• Maple Supplement to Nagle/Saff, Addison–Wesley, 1996 Meade

• Maple V for Engineers, Addison–Wesley, 1997 Meade and Bourkoff, (a module the Engineer’s Toolkit series)